OpenStudy (anonymous):

what is the y-coordinate of the graphs of 2x minus 3y = -1 and 2x + y = 67

OpenStudy (anonymous):

2x - 3y = -1 2x + y = 67 Subtract the two equations: -4y = -68 y = 17

OpenStudy (anonymous):

wait what do you mean subtract the two equations?

OpenStudy (anonymous):

(2x-3y)-(2x+y) = (1)-(67)

OpenStudy (anonymous):

how do you get -68?

OpenStudy (anonymous):

-1 - 67 = -68

OpenStudy (anonymous):

oh I get it now... thank you

OpenStudy (anonymous):

can you help me on another problem.... Simplify x negative cubed times y to the 4th power/x negative squared times y negative cubed?